OpenStudy (anonymous):

what is the expression in factored form? 9x^2-18x+9

OpenStudy (charlotte123):

@KLC14720 9(x -1)^2 ^_^ http://www.wolframalpha.com/input/?i=9x%5E2-18x%2B9+factored :D if ya need more help on these, I will be glad to help out ^_^

OpenStudy (anonymous):

how did u get that

OpenStudy (charlotte123):

@KLC14720 9x² - 18x + 9 = 9(x² - 2x +1) = 9(x-1)² :D

OpenStudy (anonymous):

i got 9(x^2-2x+1) but i dont understand how u got 9(x-1)^2